Which of the following is NOT a root of
x(x + 3)(x + 3)(x – 1)(2x + 1) = 0?
a. 0 c. –1
b. –3 d. 1/2

Answer by Cromlix(4381)

Hi there,
x(x + 3)(x + 3)(x – 1)(2x + 1) = 0
Roots are:
x = 0
......
x + 3 = 0
x = -3
......
x - 1 = 0
x = 1
......
2x + 1 = 0
2x = -1
x = -1/2
......
Roots are:
0, -3, 1 and -1/2
The answer would be (c).
Hope this helps:-)